# Multi-Modal Language Models as Text-to-Image Model Evaluators

### v1

@@ Appendix A MT2IE For Aesthetic Quality Evaluations
To demonstrate MT2IE’s ability to generalize beyond evaluating image-prompt consistency, we use it to evaluate image aesthetic quality and compare its aesthetic quality judgments to a state-of-the-art aesthetics scoring model. MT2IE is easily adaptable to be used to evaluate other aspects of generated images by simply using a different set of MLLM system prompts; for aesthetic quality evaluation we use the system prompts in Figure figure 8. All other aspects of MT2IE are kept the same, and we use the 800 generated prompts and images from Section section 5. We compare aesthetic scores produced by MT2IE and the aesthetics predictor of LAION (2022), a popular aesthetics-scoring model (Kirstain et al., 2023; Wu et al., 2023b; Xu et al., 2023). MT2IE ’s ranking of 8 T2I models by average aesthetic score closely matches the pre-trained aesthetics-scoring model’s ranking, with a Kendall’s $\tau$ of $0.691$. Per-image aesthetic scores of 800 images are also positively correlated with a Kendall’s $\tau$ of $0.538$. These results, in tandem with our main paper experiments, show that MT2IE can effectively evaluate aesthetic quality and prompt consistency. Additionally, we demonstrate that MT2IE can be easily repurposed to evaluate axes other than prompt consistency by rewriting the MLLM prompt used for judgement.
@@ Appendix B Ease of Adoption
The computational cost of one evaluation run with MT2IE is 40 inference passes of the utilized MLLM: 20 prompt generations with one prompt-image judgment each. Exact time and memory costs depend on MLLM size. However MT2IE greatly reduces the computational cost of generating images, as it is capable of matching T2I model rankings when evaluating on 20x less image-text pairs, which offsets the MLLM computation costs. MLLM system prompts for MT2IEare provided in Appendix appendix E and appendix C, and can be easily used with any off-the-shelf MLLM. Our code should be released upon acceptance.
@@ Appendix C VQA Accuracy Scoring Results and MLLM System Prompts
In Section section 4, we validate that MLLMs are able to effectively score T2I models. We have two scoring paradigms: VQA Accuracy (question generation and answering) and VQAScore. VQA Accuracy requires generating multiple choice questions given a description. In the main paper, we only discuss VQAScore prompt alignment scoring as it acheived the highest correlation with human judgment in experiments in Section section 4. We therefore provide more details on the VQA Accuracy process below.
We provide the prompt given to each MLLM for question generation and validation (we only use questions for which the answer to the validation prompt is ”yes”). Figure 12 shows Llama-3.2-11B-Vision Instruction, Figure 13 shows Molmo-7b-d and Figure 14 shows Llava-v1.6-34b. We also provide experimental results for MT2IEwith both prompt alignment scoring paradigms discussed in Figures figure 10 and figure 11.
@@
Figure 10: Average rank of 8 T2I models (shown on x axis where 1=highest scoring, 8=lowest scoring) evaluated by MT2IEwith adaptive prompting and progressively difficult prompt generation, where each prompt generation method is run with both question generation and answering (VQA Accuracy) or VQAScore for prompt alignment scoring. All MT2IEsettings evaluate T2I models over 20 generated prompts, and we compare to ranks from GenAIBench with CLIP-FlanT5 VQAScore. MT2IErankings generally match GenAIBench’s T2I model rankings, while the Downsampled GenAIBench model rankings are uninformative as all models have roughly the same mean rank over all seeds and high variation. See Figure figure 4 for details on GenAIBench and Downsampled GenAIBench.
Figure 11: We evaluate 8 T2I models, comparing MT2IEwith progressively difficult prompt generation, and scoring prompt alignment with both question generation and answering (VQA Accuracy) or VQAScore, to other evaluation approaches. Scores from MT2IE are obtained from just 20 prompts that are both generated and scored by an MLLM. Scores from GenAIBench are obtained by generating and scoring 1600 human-written prompts. We find consistent T2I model rankings and scores between MT2IE, using only 20 prompts, and the full GenAIBench with both VQA Accuracy and VQAScore prompt alignment scoring methods (discussed in section 4). Just downsampling GenAIBench to 20 prompts breaks the relative model ranking.
@@ Appendix D Analysis of Prompt Difficulty Metrics
For MT2IEwith Adaptive Prompting, discussed in Section section 6, we multiply prompt alignment scores with a prompt difficulty metric, in order to standardize prompt alignment scores since all models are no longer evaluated on the same prompts. An ideal prompt difficulty metric reflects how difficult it is for a T2I model to generate an image faithful to the prompt, and can weight the final prompt alignment score appropriately: scaling difficult prompts to higher values and easy prompts to lower values.
However to our knowledge, there is no existing work exploring metrics for T2I prompt difficulty. So we considered various NLP metrics for language complexity as prompt difficulty proxy metrics, including: word count, syllable count, Yngve score (Yngve, 1960) (a measure of the depth of a syntactic parse tree), Flesch–Kincaid scores (Flesch, 1948), and perplexity (Jelinek et al., 1977). We did preliminary analysis on a small set of MLLM-generated prompts and VQAScores and found that Yngve score was the metric most correlated with VQAScore, so we used Yngve score as the prompt difficulty weight in Section section 6’s experiments. After conducting all our experiments, we now have 20,640 (MLLM-generated) prompts along with their corresponding T2I model VQAScores, to analyze in order to identify a potential metric for prompt difficulty.
@@
Table table 4 shows correlations between various metrics calculated on prompts and T2I model VQAScores for those same prompts. Interestingly, simple statistics such as average word length and syllable count perform the best, outperforming semantic complexity metrics like Flesh-Kincaid grade level or Yngve score. Effectively measuring prompt difficulty is an interesting area of future work that we believe is very useful to the field of T2I evaluation.
@@ Appendix E MLLM System Prompts for Prompt Generation
MT2IE Iterative Difficulty Prompt Generation (Section section 5)
For prompt generation in Section section 5, we use the system prompt shown in Figure figure 15 and run generation with temperature 0.3 and no top-k sampling.
@@
MT2IE Adaptive Prompting Prompt Generation (Section section 6)
For adaptive prompt generation in Section section 6, we use the system prompts shown in Figures figure 16, figure 17, figure 18 and run generation with temperature 0.3 and no top-k sampling.
@@ Appendix F Ablations on Prompt Iterations
MT2IEgenerates a configurable number of prompts: in the Iterative Difficulty setting discussed in Section section 5 pronpts are progressively more difficulty in each iteration, whereas in the Adaptive Prompting setting discussed in Section section 6 each iteration’s prompt is rewritten by the MLLM based on the previous prompt’s alignment score.
Rankings of 8 T2I models for prompt iterations={2, 5, 10, 12} for Iterative Difficulty are shown in Figure figure 19 and Adaptive Prompting in Figure figure 20. Additionally we calculate Kendall’s $\tau$ rank correlation coefficient for all T2I model rankings, using the ranking obtained from GenAIBench as the ground truth ranking, shown in Table table 5. We found that using 5 prompt iterations, i.e.,  20 total prompts, with MT2IEyields strongly-correlated model scores with the GenAIBench benchmark which contains 1600 prompts. 5 prompt iterations results in the highest ranking correlation and smaller standard deviations in model rankings. We suspect that using more prompt iterations in the Iterative Difficulty setting can result in prompts that are too hard for many T2I models, resulting in uninformative scores and model rankings.
@@
Overall, TIFAv1.0 has an average of 6.87 questions per prompt and Llava-Gen TIFA has an average of 6.14 questions per prompt. This results in Llava-Gen TIFA containing around 10% less questions than TIFAv1.0 while covering the same number of prompts. Despite this, Llava-Gen TIFA achieves higher correlations with human judgment than TIFAv1.0, supporting our claims that MLLM-generated T2I benchmarks are more efficient. The distribution of question counts per prompt are shown in Figure figure 21.
@@
Across all prompts, TIFAv1.0 has an average of 4.68 elements per prompt and Llava-Gen TIFA has 4.13 elements per prompt, indicating an extremely similar number of elements identified for each prompt. For prompts with mismatched elements, the average number of mismatched elements is only 1.13, and we show counts of prompts with mismatched elements in Figure figure 22. The average BLEU (Papineni et al., 2002) score between TIFAv1.0 and Llava-Gen TIFA questions about the same elements is 0.65, which indicates a high level of similarity between questions covering the same concepts.

### v2

@@ Appendix A MT2IE For Aesthetic Quality Evaluations
To demonstrate MT2IE’s ability to generalize beyond evaluating image-prompt consistency, we use it to evaluate image aesthetic quality and compare its aesthetic quality judgments to a state-of-the-art aesthetics scoring model. MT2IE is easily adaptable to be used to evaluate other aspects of generated images by simply using a different set of MLLM system prompts; for aesthetic quality evaluation we use the system prompts in Figure 8. All other aspects of MT2IE are kept the same, and we use the 800 generated prompts and images from Section 5. We compare aesthetic scores produced by MT2IE and the aesthetics predictor of LAION (2022), a popular aesthetics-scoring model (Kirstain et al., 2023; Wu et al., 2023b; Xu et al., 2023). MT2IE ’s ranking of 8 T2I models by average aesthetic score closely matches the pre-trained aesthetics-scoring model’s ranking, with a Kendall’s $\tau$ of $0.691$. Per-image aesthetic scores of 800 images are also positively correlated with a Kendall’s $\tau$ of $0.538$. These results, in tandem with our main paper experiments, show that MT2IE can effectively evaluate aesthetic quality and prompt consistency. Additionally, we demonstrate that MT2IE can be easily repurposed to evaluate axes other than prompt consistency by rewriting the MLLM prompt used for judgement.
@@ Appendix B Ease of Adoption
The computational cost of one evaluation run with MT2IE is 40 inference passes of the utilized MLLM: 20 prompt generations with one prompt-image judgment each. Exact time and memory costs depend on MLLM size. However MT2IE greatly reduces the computational cost of generating images, as it is capable of matching T2I model rankings when evaluating on 20x less image-text pairs, which offsets the MLLM computation costs. MLLM system prompts for MT2IEare provided in Appendix E and C, and can be easily used with any off-the-shelf MLLM. Our code should be released upon acceptance.
@@ Appendix C VQA Accuracy Scoring Results and MLLM System Prompts
In Section 4, we validate that MLLMs are able to effectively score T2I models. We have two scoring paradigms: VQA Accuracy (question generation and answering) and VQAScore. VQA Accuracy requires generating multiple choice questions given a description. In the main paper, we only discuss VQAScore prompt alignment scoring as it acheived the highest correlation with human judgment in experiments in Section 4. We therefore provide more details on the VQA Accuracy process below.
We provide the prompt given to each MLLM for question generation and validation (we only use questions for which the answer to the validation prompt is ”yes”). 12 shows Llama-3.2-11B-Vision Instruction, 13 shows Molmo-7b-d and 14 shows Llava-v1.6-34b. We also provide experimental results for MT2IEwith both prompt alignment scoring paradigms discussed in Figures 10 and 11.
@@
Figure 10: Average rank of 8 T2I models (shown on x axis where 1=highest scoring, 8=lowest scoring) evaluated by MT2IEwith adaptive prompting and progressively difficult prompt generation, where each prompt generation method is run with both question generation and answering (VQA Accuracy) or VQAScore for prompt alignment scoring. All MT2IEsettings evaluate T2I models over 20 generated prompts, and we compare to ranks from GenAIBench with CLIP-FlanT5 VQAScore. MT2IErankings generally match GenAIBench’s T2I model rankings, while the Downsampled GenAIBench model rankings are uninformative as all models have roughly the same mean rank over all seeds and high variation. See Figure 4 for details on GenAIBench and Downsampled GenAIBench.
Figure 11: We evaluate 8 T2I models, comparing MT2IEwith progressively difficult prompt generation, and scoring prompt alignment with both question generation and answering (VQA Accuracy) or VQAScore, to other evaluation approaches. Scores from MT2IE are obtained from just 20 prompts that are both generated and scored by an MLLM. Scores from GenAIBench are obtained by generating and scoring 1600 human-written prompts. We find consistent T2I model rankings and scores between MT2IE, using only 20 prompts, and the full GenAIBench with both VQA Accuracy and VQAScore prompt alignment scoring methods (discussed in 4). Just downsampling GenAIBench to 20 prompts breaks the relative model ranking.
@@ Appendix D Analysis of Prompt Difficulty Metrics
For MT2IEwith Adaptive Prompting, discussed in Section 6, we multiply prompt alignment scores with a prompt difficulty metric, in order to standardize prompt alignment scores since all models are no longer evaluated on the same prompts. An ideal prompt difficulty metric reflects how difficult it is for a T2I model to generate an image faithful to the prompt, and can weight the final prompt alignment score appropriately: scaling difficult prompts to higher values and easy prompts to lower values.
However to our knowledge, there is no existing work exploring metrics for T2I prompt difficulty. So we considered various NLP metrics for language complexity as prompt difficulty proxy metrics, including: word count, syllable count, Yngve score (Yngve, 1960) (a measure of the depth of a syntactic parse tree), Flesch–Kincaid scores (Flesch, 1948), and perplexity (Jelinek et al., 1977). We did preliminary analysis on a small set of MLLM-generated prompts and VQAScores and found that Yngve score was the metric most correlated with VQAScore, so we used Yngve score as the prompt difficulty weight in Section 6’s experiments. After conducting all our experiments, we now have 20,640 (MLLM-generated) prompts along with their corresponding T2I model VQAScores, to analyze in order to identify a potential metric for prompt difficulty.
@@
Table 4 shows correlations between various metrics calculated on prompts and T2I model VQAScores for those same prompts. Interestingly, simple statistics such as average word length and syllable count perform the best, outperforming semantic complexity metrics like Flesh-Kincaid grade level or Yngve score. Effectively measuring prompt difficulty is an interesting area of future work that we believe is very useful to the field of T2I evaluation.
@@ Appendix E MLLM System Prompts for Prompt Generation
MT2IE Iterative Difficulty Prompt Generation (Section 5)
For prompt generation in Section 5, we use the system prompt shown in Figure 15 and run generation with temperature 0.3 and no top-k sampling.
@@
MT2IE Adaptive Prompting Prompt Generation (Section 6)
For adaptive prompt generation in Section 6, we use the system prompts shown in Figures 16, 17, 18 and run generation with temperature 0.3 and no top-k sampling.
@@ Appendix F Ablations on Prompt Iterations
MT2IEgenerates a configurable number of prompts: in the Iterative Difficulty setting discussed in Section 5 pronpts are progressively more difficulty in each iteration, whereas in the Adaptive Prompting setting discussed in Section 6 each iteration’s prompt is rewritten by the MLLM based on the previous prompt’s alignment score.
Rankings of 8 T2I models for prompt iterations={2, 5, 10, 12} for Iterative Difficulty are shown in Figure 19 and Adaptive Prompting in Figure 20. Additionally we calculate Kendall’s $\tau$ rank correlation coefficient for all T2I model rankings, using the ranking obtained from GenAIBench as the ground truth ranking, shown in Table 5. We found that using 5 prompt iterations, i.e.,  20 total prompts, with MT2IEyields strongly-correlated model scores with the GenAIBench benchmark which contains 1600 prompts. 5 prompt iterations results in the highest ranking correlation and smaller standard deviations in model rankings. We suspect that using more prompt iterations in the Iterative Difficulty setting can result in prompts that are too hard for many T2I models, resulting in uninformative scores and model rankings.
@@
Overall, TIFAv1.0 has an average of 6.87 questions per prompt and Llava-Gen TIFA has an average of 6.14 questions per prompt. This results in Llava-Gen TIFA containing around 10% less questions than TIFAv1.0 while covering the same number of prompts. Despite this, Llava-Gen TIFA achieves higher correlations with human judgment than TIFAv1.0, supporting our claims that MLLM-generated T2I benchmarks are more efficient. The distribution of question counts per prompt are shown in Figure 21.
@@
Across all prompts, TIFAv1.0 has an average of 4.68 elements per prompt and Llava-Gen TIFA has 4.13 elements per prompt, indicating an extremely similar number of elements identified for each prompt. For prompts with mismatched elements, the average number of mismatched elements is only 1.13, and we show counts of prompts with mismatched elements in Figure 22. The average BLEU (Papineni et al., 2002) score between TIFAv1.0 and Llava-Gen TIFA questions about the same elements is 0.65, which indicates a high level of similarity between questions covering the same concepts.